# Inhibition of splitting of the chiral and deconfinement transition due to rotation in QCD: the phase diagram of linear sigma model coupled to Polyakov loop

### v1

@@ III.1 Co-rotating frame
$e_{\hat{t}}$ $=\partial_{t}-\Omega(y\partial_{x}-x\partial_{y})\,,\qquad$ $\omega^{\hat{t}}=dt\,,$
$e_{\hat{x}}$ $=\partial_{x}\,,\qquad$ $\omega^{\hat{x}}=dx+\Omega ydt\,,$ (32)
$e_{\hat{y}}$ $=\partial_{y}\,,\qquad$ $\omega^{\hat{y}}=dy-\Omega xdt\,,$
@@
The connection coefficients $\Gamma_{\hat{\rho}\hat{\sigma}\hat{\alpha}}=\frac{1}{2}(c_{\hat{\rho}\hat{\sigma}\hat{\alpha}}+c_{\hat{\rho}\hat{\sigma}\hat{\alpha}}-c_{\hat{\rho}\hat{\sigma}\hat{\alpha}})$ vanish identically, except for the following component:
$\Gamma_{\hat{x}\hat{y}\hat{t}}=-\Omega\,.$ (35)
@@
With the above convention, the single non-vanishing spin connection coefficient takes the simple form, $\Gamma_{\hat{t}}=i\Omega S^{z}$, implying
@@
are the operators of the total angular momentum and its spin part, respectively.
@@
Below, the thermodynamic quantities will be evaluated in the Euclidean spacetime in the Matsubara formalism. Under the Wick rotation to the imaginary time, $t\to\tau=it$, the partition function of the model gets the form:
@@ III.2 Dirac modes and spectral boundaries
where we denote the components of the background field $(A_{0;1},{\mathcal{A}}_{0;2},{\mathcal{A}}_{0;3})=(\phi,\phi^{\prime},-\phi-\phi^{\prime})$,
@@
$L_{+}$ $\equiv L=\frac{1}{3}\bigl{(}e^{i\phi}+e^{i\phi^{\prime}}+e^{-i\phi-i\phi^{\prime}}\bigr{)}\,,$ (50)
$L_{-}$ $\equiv L^{*}=\frac{1}{3}\bigl{(}e^{-i\phi}+e^{-i\phi^{\prime}}+e^{i\phi+i\phi^{\prime}}\bigr{)}\,.$ (51)
The modes $U_{j}$ must satisfy boundary conditions on the enclosing cylinder of radius $R$. In this paper, we employ the spectral boundary conditions Hortacsu et al. (1980) as discussed for the cylindrical setup in Ref. Ambrus and Winstanley (2016). These conditions provide necessary and sufficient conditions to yield a consistent quantization by requiring the self-adjointness of the Hamiltonian. In terms of the eigenspinors $U_{j}$, the spectral condition amounts to the vanishing of either the top and third (for $m>0$) or the second and fourth components of the spinor $U_{j}$ (for $m<0$). These conditions guarantee the conservation of the total vector (baryon) charge inside the cylinder, as well as the self-adjointness of the Hamiltonian.
@@ III.3 Free energy of the rotating system
$=\ln[1+L_{\varsigma}e^{-\beta\widetilde{\mathcal{E}}_{\varsigma}}+L_{-\varsigma}e^{-2\beta\widetilde{\mathcal{E}}_{\varsigma}}+e^{-3\beta\widetilde{\mathcal{E}}_{\varsigma}}]\,.$ (57)
@@
So far, in this section, we have only discussed the thermodynamics of the fermionic part of the model. The bosonic part is encoded in the classical condensate $\sigma$, which is not affected by rotation directly. The effect of rotation on the field $\sigma$ appears only through the interaction with fermionic loops, which do feel the rotation directly, as we discussed in this section. Moreover, we treat the condensate $\sigma$ in the homogeneous approximation that is used in most of the studies so far. Namely, the condensate is set to be a coordinate-independent quantity. While noticing that the homogeneous approximation allows us to compare our results with the results of other approaches, we acknowledge that the realistic rotating plasma should develop a radial inhomogeneity, which should become rather pronounced for fast rotation Chernodub (2021); Chernodub et al. (2023); Braguta et al. (2024b).
@@ IV Splitting of transitions
In this section, we thoroughly evaluate the phase diagram of the PLSMq inside the cylindrical volume following our strategy described in Section III.
@@ IV.1.1 Softening, shifting, and splitting of transitions
First, we consider a non-rotating case with $\Omega=0$. In Fig. 4, we show the mean values of the chiral condensate and the Polyakov loops as the functions of temperature for various radii $R$ of the cylinder at chemical potentials $\mu=0$ and $\mu=0.25$ GeV. In an infinite volume and for the vanishing chemical potential, $\mu=0$, the transition has the nature of the smooth crossover since no order parameters show anything similar to a critical behavior. At $\mu=0.25$ GeV, the $R\to\infty$ transition is a strong first-order phase transition. What happens with these transitions when we decrease the radius of the cylinder $R$?
@@
The described softening of the transition temperature is most clearly seen for the condensate $\sigma$ for the large chemical potential $\mu=0.25$ GeV. As the radius diminishes, the jump in the condensate diminishes and, at $R=2$ fm, is not seen at all. This radius corresponds to the energy scale in 100 MeV, which is compatible with the pion mass. Therefore, the restriction of one of the dimensions to this value leads to a drastic softening effect on the phase transition.
Second, we find that the decrease in the radius of the cylinder $R$ leads to an increase in the pseudocritical temperatures of both crossovers. This effect can also be anticipated given the fact that the decrease of the radius leads to an enhancement of the thermal contribution to the chiral condensate inside the cylinder Ambrus and Winstanley (2016). In order to break a stronger chiral condensate, higher temperatures are needed. Due to the coupling between the chiral and confining degrees of freedom in the PLSMq, the expectation value of the Polyakov loop follows the same behavior. The enhancement of the critical temperature is, therefore, a finite-volume phenomenon.
@@ IV.1.2 Quantifying the splitting of transition temperatures
How strong is the splitting in chiral and deconfining temperatures? To this end, we found the inflection points in the chiral order parameter, $\sigma$, and in the deconfining order parameter, $L$ (determined, as usual, along a diagonal line crossing the origin of the parameter space, $T=\mu=0$). We plot the crossover temperatures in Fig. 5(a) at a fixed radius of the system $R$ as a function of the chemical potential $\mu$ and vice versa, at a set of fixed chemical potentials as the function of the radius of the cylinder $R$.
@@
Figure 5(b) allows us to quantify the magnitude of the splitting as a function of the radius. For presentation reasons, it is convenient to plot the split in the pseudo-critical temperatures of chiral, $T_{c}^{(\sigma)}$, and deconfining, $T_{c}^{(L)}$, transitions:
@@ IV.2 Effect of rotation on temperature splitting
Figure 6: Effect of rotation on the separation of chiral and deconfinement crossovers: (a) the pseudocritical temperatures as a function of $R$ and (b) their difference $\Delta T_{c}$ (59) as a function of $1/R$ at different values of $\Omega R$. The results are shown for vanishing chemical potential $\mu=0$, which gives the maximum temperature difference $\Delta T_{c}$. In panel (a), the empty (filled) symbols correspond to chiral (deconfinement) transitions.
Let us relate the results shown in Fig. 6 to the physical conditions corresponding to the realistic conditions of the vortical plasmas produced in the relativistic heavy-ion collisions. In the RHIC experiment, the vorticity of the quark-gluon plasma corresponds to the angular frequency of $\Omega\simeq 6.6$ MeV Adamczyk et al. (2017), which translates to $\Omega R\simeq 0.165$ for a system of a rather large size $R=5$ fm. Thus, the rotation velocity at the boundary of the system is much smaller than the ultrarelativistic limit $\Omega R=1$. To put this number in a different perspective, the rotational Lorentz factor at the edge of the system for RHIC-type collisions amounts to $1/\sqrt{1-(\Omega R)^{2}}\approx 1.01$, providing us with a minuscule one-percent increase. In other words, the effect of the presence of the boundary of a rotating system on the thermodynamic properties of this system is much larger than the effect of the rotation itself.
@@
Figure 7: Difference in the crossover temperatures $\Delta T_{c}$ for chiral and confinement-deconfinement transitions at different values of the chemical potential $\mu$ as a function of angular velocity, given in units of $\Omega R$. Each plot corresponds to a fixed radius of the cylinder, $R=(1,2,3,5)$ fm.
@@
Figure 8: Variation of (a) the $\sigma$ and $L$ order parameters; and (b) the slopes $\partial\frac{\sigma}{f_{\pi}}/\partial T$ and $\partial L/\partial T$, with respect to the temperature $T$, for a system of size $R=3$ fm, at various values of $\Omega R$ at $\mu=0$.
@@
Finally, we would like to comment on the nature of the crossover transition in the regime where the chiral and deconfinement transition temperatures do not coincide. Panel (a) of Fig. 8 shows the temperature dependence of both the sigma meson $\sigma$ and the Polyakov loop $L$, in the case when $R=3$ fm and $\mu=0$, for $\Omega R=0$, $0.8$ and $1$. These values of $\Omega R$ were chosen to correspond to the maximum value of $\Delta T_{pc}$, its minimum value, and again the peak reached at the causality limit, respectively. Looking at the plot, we can see that the chiral and deconfinement transitions are very much overlapped and no clear distinction between their respective pseudocritical transition temperatures can be seen. Panel (b) of the same figure shows the scaled slopes $\mathcal{N}\partial\sigma/\partial T$ and $\partial L/\partial T$, in the above cases. Here the normalization $\mathcal{N}$ is such that the peak of $\partial\sigma/\partial T$ lies at the same value as that of $\partial L/\partial L$. While the peaks of these slopes, used by us to identify the corresponding pseudocritical temperature, do not coincide, the bell-shaped area developing around these peaks is overlapped. We can thus conclude that a discussion of a unique value for the pseudocritical temperature, and hence the identification of a split between deconfinement and chiral symmetry restoration, is at best ambiguous in the case of rotating systems.
@@ V Conclusions
We have found that the confinement and chiral transitions split in the system. However, this splitting effect originates not due to the rotation of the system but rather due to its finite volume. In a very small non-rotating system with a radius of one fermi, the splitting can reach a modest $\Delta T_{c}\sim 10$ MeV, which is comparable with the width of the crossover transition Aoki et al. (2006); Borsanyi et al. (2010); Bazavov et al. (2012). For systems of the size of about 3-5 fm, the splitting becomes very negligible, $\Delta T\sim 1$ MeV.
@@
One should stress that we used the uniform approximation, in which both the chiral condensate and the expectation value of the Polyakov loop are treated as coordinate-independent quantities. This approach, which can be justified for a central region of plasma that rotates with a slow rotational velocity, $\Omega R\ll 1$, can not be applied for rapidly rotating, in the sense of the boundary velocity, systems with $\Omega R\sim 1$. For such plasmas, the inhomogeneities must be taken into account. Still, we believe that the results of this paper, obtained in the uniform approximation, are applicable to realistic plasmas that rotate relatively slowly.

### v2

@@ III.1 Co-rotating frame
$e_{\hat{t}}$ $=\partial_{t}+\Omega(y\partial_{x}-x\partial_{y})\,,\qquad$ $\omega^{\hat{t}}=dt\,,$
$e_{\hat{x}}$ $=\partial_{x}\,,\qquad$ $\omega^{\hat{x}}=dx-\Omega ydt\,,$ (32)
$e_{\hat{y}}$ $=\partial_{y}\,,\qquad$ $\omega^{\hat{y}}=dy+\Omega xdt\,,$
@@
The connection coefficients $\Gamma_{\hat{\rho}\hat{\sigma}\hat{\alpha}}=\frac{1}{2}(c_{\hat{\rho}\hat{\sigma}\hat{\alpha}}+c_{\hat{\rho}\hat{\alpha}\hat{\sigma}}-c_{\hat{\sigma}\hat{\alpha}\hat{\rho}})$ vanish identically, except for the following component:
$\Gamma_{\hat{x}\hat{y}\hat{t}}=\Omega\,.$ (35)
@@
With the above convention, the single non-vanishing spin connection coefficient takes the simple form, $\Gamma_{\hat{t}}=-i\Omega S^{z}$, implying
@@
are the operator of the total angular momentum and its spin part, respectively.
@@
Below, the thermodynamic quantities will be evaluated in the Euclidean spacetime in the Matsubara formalism. Under the Wick rotation to the imaginary time, $t\to\tau=it$, the partition function of the model takes the form:
@@ III.2 Dirac modes and spectral boundaries
where we denote the components of the background field $(\mathcal{A}_{0;1},{\mathcal{A}}_{0;2},{\mathcal{A}}_{0;3})=(\phi,\phi^{\prime},-\phi-\phi^{\prime})$,
@@
$L_{+}$ $\equiv L=\frac{1}{3}\bigl{(}e^{\beta\phi}+e^{\beta\phi^{\prime}}+e^{-\beta(\phi+\phi^{\prime})}\bigr{)}\,,$ (50)
$L_{-}$ $\equiv L^{*}=\frac{1}{3}\bigl{(}e^{-\beta\phi}+e^{-\beta\phi^{\prime}}+e^{\beta(\phi+\phi^{\prime})}\bigr{)}\,.$ (51)
The modes $U_{j}$ must satisfy boundary conditions on the enclosing cylinder of radius $R$. In this paper, we employ the spectral boundary conditions Hortacsu et al. (1980) as discussed for the cylindrical setup in Ref. Ambrus and Winstanley (2016). These conditions provide necessary and sufficient conditions to yield a consistent quantization by imposing the self-adjointness of the Hamiltonian. In terms of the eigenspinors $U_{j}$, the spectral condition amounts to the vanishing of either the top and third (for $m_{j}>0$) or the second and fourth components of the spinor $U_{j}$ (for $m_{j}<0$). These conditions guarantee the conservation of the total vector (baryon) charge inside the cylinder, as well as the self-adjointness of the Hamiltonian.
@@ III.3 Free energy of the rotating system
$=\ln[1+3L_{\varsigma}e^{-\beta\widetilde{\mathcal{E}}_{\varsigma}}+3L_{-\varsigma}e^{-2\beta\widetilde{\mathcal{E}}_{\varsigma}}+e^{-3\beta\widetilde{\mathcal{E}}_{\varsigma}}]\,.$ (57)
@@
So far, in this section, we have only discussed the thermodynamics of the fermionic part of the model. The bosonic part is encoded in the classical condensate $\sigma$, which is not affected by rotation directly. The effect of rotation on the field $\sigma$ appears only through the interaction with fermionic loops, which do feel the rotation directly, as we discussed in this section. Moreover, we treat the condensate $\sigma$ in the homogeneous approximation that is used in most of the studies so far. Namely, the condensate is set to be a coordinate-independent quantity. While noticing that the homogeneous approximation allows us to compare our results with the results of other approaches, we acknowledge that the realistic rotating plasma should develop a radial inhomogeneity, which should become rather pronounced for fast rotation Chernodub (2021); Chernodub et al. (2023); Braguta et al. (2024b). We will address the inhomogeneous thermodynamic ground state of the vortical plasma elsewhere.
@@ IV Splitting of transitions
In this section, we thoroughly evaluate the phase diagram of the PLSMq inside the cylindrical volume following our strategy described in Section III. While the static PLSM has been intensively investigated in the literature, the effects of the cylindrical boundaries on the phase diagram have never been studied in this model. Noticing that the boundary conditions are essential for our study to keep causality in place, we investigate the effect of boundaries on the splitting of the chiral and deconfinement transitions below.
@@ IV.1.1 Softening, shifting, and splitting of transitions
First, we consider the non-rotating case with $\Omega=0$. In Fig. 4, we show the mean values of the chiral condensate and the Polyakov loops as functions of temperature for various radii $R$ of the cylinder at chemical potentials $\mu=0$ and $\mu=0.25$ GeV. In an infinite volume and for vanishing chemical potential, $\mu=0$, the transition has the nature of a smooth crossover, since none of the order parameters show anything similar to critical behavior. At $\mu=0.25$ GeV, the $R\to\infty$ transition is a strong first-order phase transition. What happens with these transitions when we decrease the radius of the cylinder $R$?
@@
The described softening of the transition temperature is most clearly seen for the condensate $\sigma$ for the large chemical potential $\mu=0.25$ GeV. As the radius diminishes, the jump in the condensate diminishes and, at $R=2$ fm, it is not seen at all. This radius corresponds to the energy scale of 100 MeV, which is compatible with the pion mass. Therefore, the restriction of one of the dimensions to this value leads to a drastic softening effect on the phase transition.
Second, we find that the decrease in the radius of the cylinder $R$ leads to an increase in the pseudocritical temperatures of both crossovers. This effect can also be anticipated given the fact that the decrease of the radius leads to a suppression of the thermal contribution to the chiral condensate inside the cylinder Ambrus and Winstanley (2016), due to an enhancement of the mass gap induced by the transverse momentum quantization. In order to overcome this gap, higher temperatures are needed. Due to the coupling between the chiral and confining degrees of freedom in the PLSMq, the expectation value of the Polyakov loop follows the same behavior. The enhancement of the critical temperature is, therefore, a finite-volume phenomenon.
@@ IV.1.2 Quantifying the splitting of transition temperatures
How strong is the splitting in chiral and deconfining temperatures? To this end, we found the inflection points in the chiral order parameter, $\sigma$, and in the deconfining order parameter, $L$. We plot these crossover temperatures in Fig. 5(a) at a fixed radius of the system $R$ determined along a diagonal line crossing the origin of the parameter space, $T=\mu=0$.
@@
Figure 5(b) allows us to quantify the magnitude of the splitting as a function of the radius. Here we evaluate the crossover temperatures along the line of constant $\mu$ and for presentation reasons, we plot the split in the pseudo-critical temperatures of chiral, $T_{c}^{(\sigma)}$, and deconfining, $T_{c}^{(L)}$, transitions:
@@ IV.2 Effect of rotation on temperature splitting
Figure 6: Effect of rotation on the separation of chiral and deconfinement crossovers: (a) the pseudocritical temperatures as a function of $R$ and (b) their difference $\Delta T_{c}$ (59) as functions of the cylinder radius $R$ and its inverse, $1/R$, respectively, at different values of $\Omega R$. The results are shown for vanishing chemical potential $\mu=0$, which gives the maximum temperature difference $\Delta T_{c}$. In panel (a), the empty (filled) symbols correspond to chiral (deconfinement) transitions.
Let us relate the results shown in Fig. 6 to the physical conditions corresponding to the realistic conditions of the vortical plasmas produced in the relativistic heavy-ion collisions. In the RHIC experiment, the vorticity of the quark-gluon plasma corresponds to the angular frequency of $\Omega\simeq 6.6$ MeV Adamczyk et al. (2017), which translates to $\Omega R\simeq 0.165$ for a system of a rather large size $R=5$ fm. Thus, the rotation velocity at the boundary of the system is much smaller than the ultrarelativistic limit $\Omega R=1$. To put this number in a different perspective, the rotational Lorentz factor at the edge of the system for RHIC-type collisions amounts to $1/\sqrt{1-(\Omega R)^{2}}\approx 1.01$, providing us with a minuscule one-percent enhancement. In other words, the effect of the presence of the boundary of a rotating system on the thermodynamic properties of this system is much larger than the effect of the rotation itself.
@@
Figure 7: Difference in the crossover temperatures $\Delta T_{c}$ for chiral and confinement-deconfinement transitions at different values of the chemical potential $\mu$ as a function of angular velocity, given in units of $\Omega R$. Each plot corresponds to a fixed radius of the cylinder, $R=(1,2,3,5)$ fm
@@
Figure 8: Variation of (a) the $\sigma$ and $L$ order parameters; and (b) the slopes $\partial\bar{\sigma}/\partial T$ $(\bar{\sigma}\equiv\sigma/f_{\pi})$ and $\partial L/\partial T$, with respect to the temperature $T$, for a system of size $R=3$ fm, at various values of $\Omega R$ at $\mu=0$. The slope $d\bar{\sigma}/dT$ is scaled by a constant factor $\mathcal{N}$ for better visibility.
@@
Finally, we would like to comment on the nature of the crossover transition in the regime where the chiral and deconfinement transition temperatures do not coincide. Panel (a) of Fig. 8 shows the temperature dependence of both the sigma meson $\sigma$ and the Polyakov loop $L$, in the case when $R=3$ fm and $\mu=0$, for $\Omega R=0$, $0.8$ and $1$. These values of $\Omega R$ were chosen to correspond to the maximum value of $\Delta T_{pc}$, its minimum value, and again the peak reached at the causality limit, respectively. Looking at the plot, we can see that the chiral and deconfinement transitions are very much overlapped and no clear distinction between their respective pseudocritical transition temperatures can be seen. Panel (b) of the same figure shows the scaled slopes $\mathcal{N}\partial\bar{\sigma}/\partial T$ and $\partial L/\partial T$, where $\bar{\sigma}\equiv\sigma/f_{\pi}$, for the above mentioned values of $\Omega R$. Here the normalization $\mathcal{N}$ is such that the peak of $\partial\bar{\sigma}/\partial T$ lies at the same value as that of $\partial L/\partial T$. While the peaks of these slopes, used by us to identify the corresponding pseudocritical temperature, do not coincide, the bell-shaped areas developing around these peaks are overlapped. We can thus conclude that a discussion of a unique value for the pseudocritical temperature, and hence the identification of a split between deconfinement and chiral symmetry restoration, is at best ambiguous in the case of bounded rotating systems.
@@ V Conclusions
We found that the confinement and chiral transitions split in the system. However, this splitting effect originates not due to the rotation of the system but rather due to its finite volume. In a very small non-rotating system with a radius of one fermi, the splitting can reach a modest $\Delta T_{c}\sim 10$ MeV, which is comparable with the width of the crossover transition Aoki et al. (2006); Borsanyi et al. (2010); Bazavov et al. (2012). For systems of the size of about 3-5 fm, the splitting becomes very negligible, $\Delta T\sim 1$ MeV.
@@
One should stress that we used the uniform approximation, in which both the chiral condensate and the expectation value of the Polyakov loop are treated as coordinate-independent quantities. This approach, which can be justified for the central region of the plasma that rotates with a slow rotational velocity, $\Omega R\ll 1$, cannot be applied for rapidly rotating systems for which the boundary velocity reaches $\Omega R\sim 1$. For such plasmas, the inhomogeneities must be taken into account. This important question will be addressed elsewhere. Nevertheless, we believe that the results of this paper, obtained in the uniform approximation, are applicable to realistic plasmas that rotate relatively slowly.